OpenStudy (anonymous):

I'm stuck on this one algebra question which I was suppose to create, yet I don't know how to solve it. Sally drove a Jet Ski for 4 hours upstream against a water current flowing at 10 MPH. When returning back home with the water current it was a 2 hour ride. Calculate how fast Sally traveling in still water? Thanks in advance for any help! :)

OpenStudy (amistre64):

hmmm, speed = distance/time speed = current +- jetski. assuming the distance is constant, and the speeds of the river and jet are constant ....

OpenStudy (amistre64):

(c-j) = d/4 4(c-j) = d 2(c+j) = d 4(c-j) = 2(c+j) 4(10-j) = 2(10+j) solve for j?

OpenStudy (amistre64):

i seem to have deleted my (c+j) = d/2 :)

OpenStudy (amistre64):

4(10-j) = 2(10+j) 2(10-j) = 10+j 20-2j = 10+j 10 = 3j 10/3 = j

OpenStudy (amistre64):

does that make sense? and did i think it thru correctly?

OpenStudy (anonymous):

I seem to understand how you solved it but got lost in the beginning when you were setting up the equation, but I got that the speed of the Jet Ski, or the speed of it in "still" water, which is the same (from my understanding) is 30 not 10/3... when solving I figured that both traveled at different times, so I would multiply each one by their respective time traveled. 4(j-10) = 2(j+10) 4j-40 = 2j+20 2j=60 j = 30 Is this wrong? How did you set up your equation? Still a bit confused... :/

OpenStudy (amistre64):

hmmm, i do tend to get these things mixed around at times :) your set up with the jet +- current seems better.

OpenStudy (amistre64):

in fact, to get upstream he would have to be going faster than the current :)

OpenStudy (anonymous):

So just to recap, the speed of the Jet ski is 30 MPH? and thanks for your input!

OpenStudy (amistre64):

yep, thats would be correct.